OpenStudy (anonymous):

F(x)=x^2 -14 and G(x)=7-x what is (F times G)(-6)?

hero (hero):

There are various ways to go about this

hero (hero):

You can either go ahead and multiply f(x)*g(x) knowing that (f*g)(x) = f(x)*g(x) or you can just know that (f*g)(-6) = f(-6)*g(-6) and do it that way So (f*g)(-6) = f(-6)*g(-6) = ((-6)^2 - 14)(7-(-6)) = (36 - 14)(7+6) = (22)(13) = 286